Session: Mar 2009 (session 1)
Wavelength: 1.3 cm
To give feedback on this experiment, click
Medicina (Roma at Tue Mar 3 17:34:56 2009) Success :
sky clody and rainy
Shanghai (Bo Xia at Tue Mar 3 17:41:05 2009) Success :
Yebes (Pablo de Vicente at Wed Mar 4 11:25:15 2009) Success :
Rainy. Took a lot of time to transfer data via FTP.
Onsala (Michael Lindqvist at Thu Mar 5 07:17:04 2009) Minor failures :
disk2file problems. Mark5 hanged.
Effelsberg (Uwe Bach at Thu Mar 5 09:04:25 2009) Minor failures :
Not onsource from 15:57 to 16:02. Much too late on source for w3oh, because of a full 360 degree turn, but position was wrong anyhow.
Metsahovi (Guifre at Thu Mar 5 17:04:14 2009) Success :
Noto (P. Cassaro at Fri Mar 6 15:52:56 2009) Success : RECE
No problems, but only LCP data, due to a receiver problem.
Jodrell_Bank (Alastair Gunn at Tue Mar 24 11:11:29 2009) Minor failures :
The MkII telescope was used (Jb2).
Disk pack WSRT-056 failed to record at experiment start. Began recording on WSRT-010 at 12:17 UT. No other known problems.